• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Solvignon

    Log entry time 00:01:56 on April17,2007

    Entry number 199303

    keyword=Shift summary

    16:00 Starting the shift with controlled access.

    17:26 MCC ready to send beam but we are not ready.

    18:19 Ready for beam permit.

    18:44 10uA. Adjusting BPM positions

    19:00 Beam positions are now good.

    19:14 Find the beam spot on BeO target. (mark new position on monitor). Now moving the target to 3He.

    19:27 The target is ready. Ask for 95uA. Rates are fine so start production.

    00:00 End of the shift. Ran smoothly since 19:37.

    --------------------------
    List of runs:

    2546 19:37-20:06 97uA, 1.2M
    2547 20:06-20:31 97uA, 1.2M
    2548 20:31-20:57 97uA, 1.2M
    2549 20:57-21:24 97uA, 1.2M
    2550 21:25-21:52 97uA, 1.2M
    2551 21:52-22:18 97uA, 1.2M
    2552 21:18-23:04 20-97uA, 1.2M
    2553 23:05-23:35 97uA, 1.2M
    2554 23:35-00:00 97uA, 1.2M